OpenStudy (anonymous):

the output of a finished product, f(x,y) can be described by the function: f(x,y) =a(x^b)y^(1-b), where x stands for the amount of money expended for labor, y stands for the amount expended on capital, and a and b are positive constants with 0

OpenStudy (turingtest):

f(x,y) =a(x^b)y^(1-b) f(px,py)=a(px)^b(py)^(1-b) =a(x^b)y^(1-b)(p^b)p^(1-b) =a(x^b)y^(1-b)p^(1+b-b) =pa(x^b)y^(1-b)=pf(x,y)

OpenStudy (anonymous):

how did u do line 3?

OpenStudy (anonymous):

using the result of part a, to show that if the amount of money expended for labor and capital are both increased by r percent, then the output is also increased by r percent.

OpenStudy (turingtest):

for line the ^b and ^(1-b) exponents get distributed (px)^b=(p^b)(x^b) (py)^(1-b)=p^(1-b)y^(1-b) I just moved the p^b next to the p^(1-b) in the same step. Sorry if that threw you.